### v3

# A central limit theorem for random tangent fields on stratified spaces

Abstract.
Variation of empirical Fréchet means on a metric space with curvature bounded above is encoded via random fields indexed by unit tangent vectors. A central limit theorem shows these random tangent fields converge to a Gaussian such field and lays the foundation for more traditionally formulated central limit theorems in subsequent work.
Key words and phrases:
Central Limit Theorem; random tangent field; barycenter (Fréchet mean); CAT(k) space; stratified space
2020 Mathematics Subject Classification:
Primary: 60F05, 53C23, 60D05, 60B05, 49J52, 62R20, 62R07, 57N80, 58A35, 58Z05, 53C80, 62G20, 62R30, 28C99; Secondary: 58K30, 57R57, 92B10
Introduction
The classical central limit theorem (CLT) captures the distributional form of the leading order fluctuation of a sample mean around its limiting value, namely the true or population mean. After rescaling, these fluctuations typically converge to a Gaussian distribution. When the samples are drawn from a linear space, such as a flat vector space, where addition of elements is well defined, both the sample and population means are given by the classical average. These classical averages can also be identified as the minimizers of a least-squared error problem involving the empirical and population distribution, which largely explains the centrality of averages in asymptotic problems. While averages do not translate to nonlinear spaces, since geometry precludes simple analogues of arithmetic averages, the minimization formulation does translate, as it requires only a metric. Minimizers of the Fréchet function
(1) $F_{\mu}(p)=\frac{1}{2}\int_{\mathcal{M}}d(p,x)^{2}\,\mu(dx)$
of the given measure $\mu$ on a space $\mathcal{M}$, known as barycenters, play the central role of the Fréchet mean of the distribution $\mu$. Furthermore, variation of the empirical barycenter from the population barycenter ${\bar{\mu}}$ is informed by the geometry of $\mathcal{M}$ and is precisely the object, once rescaled, to be described via the CLT.
In settings where the geometric structure of $\mathcal{M}$ is smooth [BP05], spatial variation in $\mathcal{M}$ maps invertibly to the tangent space, at least locally near the population mean ${\bar{\mu}}$. This allows one to consider the limit of the scaled fluctuations in the tangent space and then push them back down to the base space $\mathcal{M}$. However, advances in geometric statistics have uncovered the need to work with singular spaces $\mathcal{M}$ (see [MMT23a], especially sources cited in the Introduction there) that may not have a smooth geometric structure. These spaces typically have singular points, whose tangent spaces are cones rather than flat, Euclidian vector spaces as in the smooth setting. In the absence of smoothness, it is impossible to transform tangent data into a linear space invertibly. Instead of taking into account all spatial variation of empirical means around ${\bar{\mu}}$, our approach here reduces the singular situation to the linear case by treating variation around the population mean radially—that is, ray by ray. The result is a potent interpretation of the CLT as convergence of real-valued random fields indexed by the unit sphere (Theorem 2). This perspective makes sense even when the underlying space $\mathcal{M}$ is neither smooth nor linear. The sequel [MMT23d] to this paper uses the results of this note on CLTs for random tangent fields to prove a more classical-looking CLT for full spatial variation of empirical means in $\mathcal{M}$ around ${\bar{\mu}}$.
To see how this proceeds, and to highlight connections with the classical case, begin by returning to a linear setting. Simplify notation by writing $F_{\mu}=F$ for the Fréchet function and $F_{n}=F_{\mu_{n}}$ for the empirical Fréchet function of the empirical measure $\mu_{n}=\frac{1}{n}\sum_{k}\delta_{X_{k}}$ corresponding to i.i.d. sample of random points $X_{1},\dots,X_{n}$ from $\mu$, so
$F_{n}(p)$ $=\frac{1}{2}\int|p-x|^{2}\mu_{n}(dx)=\frac{1}{2}\int\bigl{(}|p|^{2}-2p\cdot x+|x|^{2}\bigr{)}\mu_{n}(dx)$
$=\frac{|p|^{2}}{2}-p\cdot{\bar{\mu}}+\!\!\int\!\frac{|x|^{2}}{2}\mu(dx)-\frac{1}{n}\sum_{i=1}^{n}(p\cdot X_{i}-p\cdot{\bar{\mu}})+\frac{1}{n}\sum_{i=1}^{n}\frac{|X_{i}|^{2}}{2}-\!\!\int\!\frac{|x|^{2}}{2}\mu(dx)$
$=F(p)-{\overline{g}}_{n}(p)+C_{n},$
where
${\overline{g}}_{n}(p)=\frac{1}{n}\sum_{i=1}^{n}(p\cdot X_{i}-p\cdot{\bar{\mu}})\quad\text{and}\quad C_{n}=\frac{1}{2n}\sum_{i=1}^{n}|X_{i}|^{2}-\frac{1}{2}\int|x|^{2}\mu(dx).$
Since $C_{n}$ is a constant that is independent of $p$, depending only on $\mu_{n}$ and $\mu$, minimizing $F_{n}(p)$ coincides with minimizing $F(p)-{\overline{g}}_{n}(p)$. Hence $\sqrt{n}\,({\bar{\mu}}_{n}-{\bar{\mu}})$ becomes
$\sqrt{n}\Bigl{(}\operatorname*{argmin}_{p}F_{n}(p)-\operatorname*{argmin}_{p}F(p)\Bigr{)}=\sqrt{n}\Bigl{(}\operatorname*{argmin}_{p}\bigl{(}F(p)-{\overline{g}}_{n}(p)\bigr{)}-\operatorname*{argmin}_{p}F(p)\Bigr{)}.$
Since the setting is linear for the moment, the coordinate system can be translated so that ${\bar{\mu}}$ is the origin. Then this equation simplifies, and the usual linear CLT becomes
[…]
This phrasing shifts perspective from a spatial view—minimize the sum of square distances to the sample points—to an angular view: maximize the sum of inner products with the sample points without unduly increasing the Fréchet function of the population measure. The task of a CLT becomes comparing these two quantities asymptotically. Asymptotics of the right-hand side of (2) can be understood by first getting a handle on convergence of the average functions ${\overline{g}}_{n}$ and then passing that perturbation through the minimization problem in (2). Before expanding on that, the first step is to identify an appropriate setting where $\mathcal{M}$ is not linear.
As in the smooth case, any singular CLT naturally occurs tangentially at the Fréchet mean ${\bar{\mu}}$. For this to occur, the space $\mathcal{M}$ must possess a space of tangent data with enough structure. To that end the setting here is $\operatorname{CAT}(\kappa)$ spaces (Section 2), a rather general class of metric spaces in which each point $p\in\mathcal{M}$ has a tangent cone $T_{p}\mathcal{M}$. $\operatorname{CAT}(\kappa)$ spaces have enough structure to define an angular pairing that serves as a replacement for the inner product in a linear space.
Using a $\operatorname{CAT}(\kappa)$ space $\mathcal{M}$ and its tangent cone $T_{\bar{\mu}}\mathcal{M}$ at the Fréchet mean, the CLT (2) is recast as an expression for the distribution of the limit
(3) $\lim_{n\to\infty}\sqrt{n}\operatorname*{argmin}_{V\in T_{\bar{\mu}}\mathcal{M}}\bigl{(}F(\exp_{\bar{\mu}}V)-{\overline{g}}_{n}(V)\bigr{)}$
on $T_{\bar{\mu}}$, where now ${\overline{g}}_{n}$ takes place in the tangent cone instead of $\mathcal{M}$ itself, so $X_{i}=\log_{\bar{\mu}}x_{i}$ for $x_{i}\sim\mu$:
(4) $\begin{split}{\overline{g}}_{n}(V)&=\frac{1}{n}\sum_{i=1}^{n}\bigl{(}\langle V,X_{i}\rangle_{\bar{\mu}}-m(V)\bigr{)}\\ \quad\text{and}\quad m(V)&=\int_{\mathcal{M}}\langle V,\log_{\bar{\mu}}x\rangle_{\bar{\mu}}\,\mu(dx)=\int_{T_{\bar{\mu}}\mathcal{M}}\langle V,U\rangle_{\bar{\mu}}\,\,{\widehat{\mu}}(dU),\end{split}$
where ${\widehat{\mu}}=(\log_{\bar{\mu}})_{\sharp}\mu$ is the pushforward of the measure $\mu$ in $\mathcal{M}$ to the cone $T_{\bar{\mu}}\mathcal{M}$. The random real-valued random field ${\overline{g}}_{n}$, indexed by the tangent cone $T_{\bar{\mu}}\mathcal{M}$, and its convergence are the principal topics of this paper. Motivated by these formulas, random tangent fields are defined in Section 2. Crucially, even though $T_{\bar{\mu}}\mathcal{M}$ is not a vector space when ${\bar{\mu}}$ is a singular point, the conical geometry of $T_{\bar{\mu}}$ (Definition 1.3) has enough structure for angles and nonnegative scaling to be defined in $T_{\bar{\mu}}\mathcal{M}$. This allows the definition of an angular pairing that plays the role of inner products in (4) when the setting is singular (Definition 1.4).
Shifting to the formulation (3) has a decisive advantage. Since the tangent cone $T_{\bar{\mu}}\mathcal{M}$ need not be a vector space, classical CLTs do not apply to collections of independent random variables taking values in $T_{\bar{\mu}}\mathcal{M}$. However, classical CLTs do apply to collections of real-valued random fields indexed by the unit sphere $S_{\bar{\mu}}\mathcal{M}$ in $T_{\bar{\mu}}\mathcal{M}$, as they can be viewed as a vector space. And ${\overline{g}}_{n}$ in (4) can be viewed as a real-valued random field indexed by $S_{\bar{\mu}}\mathcal{M}$ which is the average of the $n$ independent random fields $V\mapsto\langle X_{k},V\rangle_{\bar{\mu}}-m(V)$. This change in perspective from CLT as convergence of random variables to CLT as convergence of random fields is a powerful theoretical advance that has added value even in classical cases; see [MMT23d, Remark 6.40].
In more detail, since each of the $n$ fields averaged in ${\overline{g}}_{n}$ is centered and independent, an appropriate CLT for random fields yields convergence of $\sqrt{n}\,{\overline{g}}_{n}$—again, viewed as a random field—to a Gaussian random field $G$; this is the main Theorem 2. This Gaussian random field has mean zero and covariance $\mathbb{E}[G(V)G(U)]=\Sigma(V,U)$, where
[…]
The convergence in Theorem 2 addresses what happens inside the $\operatorname*{argmin}$ from (3), but it falls short of relating the limit (3) to asymptotic spatial variation of empirical Fréchet means $\lim_{n\to\infty}\sqrt{n}\log_{\bar{\mu}}{\bar{\mu}}_{n}$ in analogy with (2). That is treated in subsequent work [MMT23d], which pushes this convergence to the tangent cone as a whole, rather than ray by ray, using tangential collapse [MMT23b] and additional theory surrounding variation of Fréchet means on a singular space $\mathcal{M}$ upon perturbation of the measure $\mu$.
Acknowledgements
DT was partially funded by DFG HU 1575/7. The other two authors were informed of [LTH23], which has related results, the day before posting this paper. JCM thanks the NSF RTG grant DMS-2038056 for general support as well as Davar Khoshnevisan and Sandra Cerrai for enlightening conversations. The referees provided extremely helpful, perceptive comments that substantively improved the paper.
1. Prerequisites
To keep this paper self-contained, recall some basic notions regarding $\operatorname{CAT}(\kappa)$ spaces and measures on them from the prequel [MMT23b]; more details, proofs, background, and examples can be found there and in the sources it cites (notably [BBI01] and [MMT23a]).
Throughout the paper, fix a complete locally compact $\operatorname{CAT}(\kappa)$ metric space $(\mathcal{M},\mathbf{d})$. Many of the definitions fix an arbitrary point ${\bar{\mu}}\in\mathcal{M}$. In practice, this point is always the Fréchet mean of a measure $\mu$ in this paper, but it does no harm to call the point ${\bar{\mu}}$ even without the presence of a measure $\mu$.
Definition 1.1.
The angle between geodesics $\gamma_{i}\colon[0,\varepsilon_{i})\to\mathcal{M}$ for $i=1,2$ emanating from ${\bar{\mu}}$ in the $\operatorname{CAT}(\kappa)$ space $(\mathcal{M},\mathbf{d})$ and parametrized by arclength is characterized by
$\cos\bigl{(}\angle(\gamma_{1},\gamma_{2})\bigr{)}=\lim_{t,s\to 0}\frac{s^{2}+t^{2}-\mathbf{d}^{2}(\gamma_{1}(s),\gamma_{2}(t))}{2st}.$
Clearly, $\angle(\gamma_{1},\gamma_{2})\in[0,\pi]$. The geodesics $\gamma_{i}$ are equivalent if the angle between them is 00. The completion $S_{\bar{\mu}}\mathcal{M}$ of the set of equivalence classes is the space of directions at ${\bar{\mu}}$.
Lemma 1.2.
The space of directions is a length space whose angular metric $\mathbf{d}_{s}$ satisfies
$\mathbf{d}_{s}(V,W)=\angle(V,W)\text{ whenever }V,W\in S_{\bar{\mu}}\mathcal{M}\text{ with }\angle(V,W)<\pi$
Proof.
This is [MMT23a, Proposition 1.7], which in turn is [BBI01, Lemma 9.1.39]. ∎
Definition 1.3.
At any point ${\bar{\mu}}$ the $\operatorname{CAT}(\kappa)$ space $\mathcal{M}$ has tangent cone
$T_{\bar{\mu}}\mathcal{M}=S_{\bar{\mu}}\mathcal{M}\times[0,\infty)/\bigl{(}S_{p}\mathcal{M}\times\{0\}\bigr{)},$
defined in the standard way as a topological quotient of the cylinder over $S_{\bar{\mu}}\mathcal{M}$, whose apex is also called ${\bar{\mu}}$. The length $\|V\|$ of a tangent vector $V$ is its $[0,\infty)$ coordinate. The unit tangent sphere of length $1$ vectors is identified with $S_{\bar{\mu}}\mathcal{M}$, the space of directions.
Definition 1.4.
Tangent vectors $V,W\in T_{\bar{\mu}}\mathcal{M}$ at any point ${\bar{\mu}}$ have angular pairing
$\langle V,W\rangle_{\bar{\mu}}=\|V\|\|W\|\cos\bigl{(}\angle(V,W)\bigr{)}.$
The subscript ${\bar{\mu}}$ is suppressed when the point ${\bar{\mu}}$ is clear from context. The conical metric on the tangent cone $T_{\bar{\mu}}\mathcal{M}$,which differs from the angular metric $d_{s}$ from above, is
$d_{\bar{\mu}}(V,W)=\sqrt{\|V\|^{2}+\|W\|^{2}-2\langle V,W\rangle_{\bar{\mu}}}\ \text{ for }V,W\in T_{\bar{\mu}}\mathcal{M}.$
Though not a true inner product, angular paring shares enough properties with an inner product, such as the Cauchy–Schwartz inequality (immediate from Lemma 1.2) and continuity with fixed basepoint ${\bar{\mu}}$ (this is [MMT23a, Lemma 1.21], which derives it from continuity of $\mathbf{d}_{s}$ in Lemma 1.2), to replace inner products in this singular setting.
Definition 1.5.
If ${\bar{\mu}}\in\mathcal{M}$ and $v$ lies in the set $\mathcal{M}^{\prime}\subseteq\mathcal{M}$ of points in the $\operatorname{CAT}(\kappa)$ space $(\mathcal{M},\mathbf{d})$ with unique shortest path to ${\bar{\mu}}$, then write $\gamma_{v}$ for the unit-speed shortest path from ${\bar{\mu}}$ to $v$ and $V$ for its direction at ${\bar{\mu}}$. The log map is defined by
$\log_{\bar{\mu}}\colon\mathcal{M}^{\prime}$ $\to T_{p}\mathcal{M}\quad\text{where}\quad v\mapsto\mathbf{d}({\bar{\mu}},v)V.$
Definition 1.6 ([MMT23b, Definition 2.1]).
[…]
  • its Fréchet mean ${\bar{\mu}}$ is unique,
  • the Fréchet function of $\mu$ is locally convex in a neighborhood of ${\bar{\mu}}$, and
  • the logarithm map $\log_{\bar{\mu}}\colon\mathcal{M}\to T_{\bar{\mu}}\mathcal{M}$ is uniquely defined $\mu$-almost surely.
The pushforward of a localized measure $\mu$ to $T_{\bar{\mu}}\mathcal{M}$ is denoted
${\widehat{\mu}}=(\log_{\bar{\mu}})_{\sharp}\mu.$
Definition 1.7.
Given a localized measure $\mu$ on a $\operatorname{CAT}(\kappa)$ space $\mathcal{M}$, the directional derivative of the Fréchet function $F$ at ${\bar{\mu}}$ is
$\nabla_{{\bar{\mu}}}F:T_{\bar{\mu}}\mathcal{M}$ $\to\mathbb{R}\quad\text{where}\quad V\mapsto\frac{d}{dt}F(\exp_{\bar{\mu}}tV)|_{t=0}$
where the exponential is a constant-speed geodesic with tangent $V$ at ${\bar{\mu}}$.
Remark 1.8.
Directional derivatives are well defined by [MMT23b, Corollary 2.7], which states that
$\nabla_{{\bar{\mu}}}F(V)=-\int_{T_{\bar{\mu}}\mathcal{M}}\langle W,V\rangle_{\bar{\mu}}\,{\widehat{\mu}}(dW).$
Definition 1.9 (Stratified $\operatorname{CAT}(\kappa)$ space).
A complete, geodesic, locally compact $\operatorname{CAT}(\kappa)$ space $(\mathcal{M},\mathbf{d})$ is stratified if it decomposes $\mathcal{M}=\bigsqcup_{j=0}^{d}\mathcal{M}^{j}$ into disjoint locally closed strata $\mathcal{M}^{j}$ so that for each $j$, the stratum $\mathcal{M}^{j}$ has closure
$\overline{\mathcal{M}^{j}}=\bigcup_{k\leq j}\mathcal{M}^{k},$
and for each stratum $\mathcal{M}^{j}$, the space $(\mathcal{M}^{j},\mathbf{d}|_{\mathcal{M}^{j}})$ is a smooth manifold with geodesic distance $\mathbf{d}|_{\mathcal{M}^{j}}$ that is the restriction of $\mathbf{d}$ to $\mathcal{M}^{j}$.
Remark 1.10.
The notion of stratified $\operatorname{CAT}(\kappa)$ space in Definition 1.9 is weaker (i.e., more inclusive) than the notion of smoothly stratified metric space from [MMT23b, Definition 3.1]: a stratified $\operatorname{CAT}(\kappa)$ space becomes a smoothly stratified metric space if its log maps are locally invertible as in [MMT23b, Definition 3.1.2]. The only location in this paper where the stratified hypothesis enters is Lemma 3.5, with the consequence that Theorem 1 assumes $\mathcal{M}$ is stratified and therefore Lemmas 5.8 and 5.9 do, as well. The rest of the arguments work for arbitrary complete locally compact $\operatorname{CAT}(\kappa)$ spaces.
2. Random tangent fields
Definition 2.1 (Random tangent field).
Fix a complete probability space $(\Omega,\mathcal{F},\mathbf{P})$ and a smoothly stratified metric space $\mathcal{M}$. A random tangent field on the unit tangent sphere $S_{\bar{\mu}}\mathcal{M}$ at a point ${\bar{\mu}}\in\mathcal{M}$ is a real-valued stochastic process $f$ indexed by $S_{\bar{\mu}}\mathcal{M}$; that is,
$f\colon\Omega\times S_{\bar{\mu}}\mathcal{M}$ $\to\mathbb{R}$
is a measurable map. Typically $\omega$ is suppressed in these notations. A centered random tangent field is such a field $f$ with expectation $\mathbb{E}f(V)=0$ for all $V\in S_{\bar{\mu}}\mathcal{M}$.
Remark 2.2.
A random tangent field $f$ can be thought of—and notated—as a collection of functions $\{f(V)\colon\Omega\to\mathbb{R}\mid V\in S_{\bar{\mu}}\mathcal{M}\}$ indexed by unit tangent vectors $V$ at ${\bar{\mu}}$. If $x=x(\omega)$ is an $\mathcal{M}$-valued random variable $\Omega\to\mathcal{M}$, and $g\colon\mathcal{M}\times S_{\bar{\mu}}\mathcal{M}\to\mathbb{R}$ is any (ordinary deterministic) function, then $g$ induces a random tangent field written
$(\omega,V)\mapsto g\bigl{(}x(\omega),V\bigr{)}.$
In the interest of brevity, we typically abuse notation to write $g(x,V)$ for this random field. For example given an $\mathcal{M}$-valued random variable $x$, the function $g\colon(y,V)\mapsto\langle\log_{\bar{\mu}}y,V\rangle_{\bar{\mu}}$ induces a random tangent field $g(x,V)$, the centered instance of which appears in Definition 2.5.
Remark 2.3.
A random tangent field $f$ can be canonically extended to a homogeneous stochastic process indexed by $T_{\bar{\mu}}\mathcal{M}$ instead of $S_{\bar{\mu}}\mathcal{M}$, where $f$ is homogeneous if $f(tV)=tf(V)$ for all real $t\geq 0$ and $V\in T_{\bar{\mu}}\mathcal{M}$. This extension is not needed in this paper, but it arises in the sequel [MMT23d, Section 3.2].
This paper studies asymptotic behavior of the centered random tangent field induced by a probability measure $\mu$ on $\mathcal{M}$. Write $F=F_{\mu}$ for the Fréchet function of $\mu$ in (1).
Definition 2.4.
For any localized probability measure $\nu$ on $\mathcal{M}$, define the
  1. 1.
    tangent mean function for $V\in S_{\bar{\mu}}\mathcal{M}$ by
    $V\mapsto m(\nu,V):=\int_{\mathcal{M}}\langle\log_{\bar{\mu}}y,V\rangle_{\bar{\mu}}\nu(dy)$
  2. 2.
    tangent covariance kernel for $(V,W)\in S_{\bar{\mu}}\mathcal{M}\times S_{\bar{\mu}}\mathcal{M}$ by
    $(V,W)\mapsto\operatorname{\Sigma}(\nu,V,W):=\int_{\mathcal{M}}\bigl{[}\langle\log_{\bar{\mu}}y,V\rangle_{\bar{\mu}}-m(\nu,V)\bigr{]}\,\bigl{[}\langle\log_{\bar{\mu}}y,W\rangle_{\bar{\mu}}-m(\nu,W)\rangle_{\bar{\mu}}\bigr{]}\,\nu(dy).$
Abusing notation again, these definitions extend to any $\mathcal{M}$-valued random variable $x\colon\Omega\to\mathcal{M}$ with law $\nu$ by setting $m(x,V)=m(\nu,V)$ and $\operatorname{\Sigma}(x,V,W)=\operatorname{\Sigma}(\nu,V,W)$, so
$m(x,V)$ $=\mathbb{E}\langle\log_{\bar{\mu}}x,V\rangle_{\bar{\mu}}$
$\text{and}\quad\operatorname{\Sigma}(x,V,W)$ $=\mathbb{E}\Bigl{[}\bigl{(}\langle\log_{\bar{\mu}}x,V\rangle_{\bar{\mu}}-m(x,V)\bigr{)}\,\bigl{(}\langle\log_{\bar{\mu}}x,W\rangle_{\bar{\mu}}-m(x,W)\bigr{)}\Bigr{]}.$
Definition 2.5.
For any $\mathcal{M}$-valued random variable $x\colon\Omega\to\mathcal{M}$, the centered random tangent field at ${\bar{\mu}}$ is
$\tau(x,V):=\langle\log_{\bar{\mu}}x,V\rangle_{\bar{\mu}}-m(x,V)\,.$
If $x$ is distributed as $\mu$ then $\tau(V)=\tau(x,V)$ is the centered tangent field induced by $\mu$.
Remark 2.6.
The centered random tangent field $\tau(x,^{\!}V)$ has mean zero and covariance $\operatorname{\Sigma}(x,^{\!}V,W)$ in that $\mathbb{E}\tau(x,^{\!}V)=0$ and $\mathbb{E}\,\tau(x,^{\!}V)\tau(x,^{\!}W)=\operatorname{\Sigma}(x,^{\!}V,W)$ for all $V,W\!\in\!S_{\bar{\mu}}\mathcal{M}$.
Remark 2.7.
Centering by subtracting the tangent mean function $m(\mu,V)$ is the same as adding the derivative $\nabla_{{\bar{\mu}}}F(V)$ by [MMT23b, Corollary 2.7] (see Remark 1.8 for the statement). In particular, the tangent mean function $m(\mu,V)=-\nabla_{{\bar{\mu}}}F(V)$ is never strictly positive (as ${\bar{\mu}}$ is a minimizer of $F$) and vanishes precisely when $\nabla_{{\bar{\mu}}}F(V)=0$. This vanishing defines the escape cone $E_{\mu}$ [MMT23b, Definition 2.12], so $m(\mu,V)\leq 0$ for all $V\in S_{\bar{\mu}}\mathcal{M}$, and $m(\mu,V)=0$ if and only if $V\in E_{\mu}$ by definition.
Definition 2.8 (Gaussian tangent field).
A Gaussian tangent field with covariance $\operatorname{\Sigma}$ is a centered random tangent field $G\colon\Omega\times S_{\bar{\mu}}\mathcal{M}\to\mathbb{R}$ such that
  1. 1.
    the covariance $(V,W)\mapsto\mathbb{E}\,G(V)G(W)$ of $G$ is $\operatorname{\Sigma}(V,W)$ for all $V,W\in S_{\bar{\mu}}\mathcal{M}$, and
  2. 2.
    for all $V_{1},\ldots,V_{n}\!\in\!S_{\bar{\mu}}\mathcal{M}$, $\bigl{(}G(V_{1}),\ldots,G(V_{n})\bigr{)}$ is multivariate Gaussian distributed.
The Gaussian tangent field induced by $\mu$ is the centered Gaussian random tangent field $G\colon\Omega\times S_{\bar{\mu}}\mathcal{M}\to\mathbb{R}$ whose covariance is $\mathbb{E}G(V)G(W)=\operatorname{\Sigma}(\mu,V,W)$.
3. Regularity of random tangent fields
The main result in this section is a Kolmogorov regularity theorem adapted to our setting of random tangent fields (Theorem 1 in §3.1) and its interpretation for Gaussians random tangent fields (Corollary 3.3 in §3.2). Presentations of these are followed by the proof of Theorem 1 (in §3.3).
3.1. A Kolmogorov regularity theorem
Theorem 1.
Fix a point ${\bar{\mu}}$ in a stratified $\operatorname{CAT}(\kappa)$ space $\mathcal{M}$. Let $f$ be a random tangent field on the unit tangent sphere $S_{\bar{\mu}}\mathcal{M}$ with
(3.1) $\mathbb{E}|f(V)-f(W)|^{p}\leq C_{p}d_{\bar{\mu}}(W,V)^{p}$
for all $p\geq 1$ for some finite constant $C_{p}$. Then there exists a version of the random tangent field $f$ which is almost surely a Hölder continuous function $S_{\bar{\mu}}\mathcal{M}\to\mathbb{R}$ for all Hölder exponents in $(0,1)$.
The proof, at the end of Section 3, and that of Corollary 3.3 need the following basic estimates.
Proposition 3.1.
Let $x$ be a random variable in $\mathcal{M}$ with $\mathbb{E}[d({\bar{\mu}},x)^{2}]<\infty$. Then for $U,V,W\in S_{\bar{\mu}}\mathcal{M}$,
$\lvert\tau(x,V)-\tau(x,U)\rvert$ $\leq\bigl{(}\mathbb{E}d({\bar{\mu}},x)+d({\bar{\mu}},x)\bigr{)}d_{\bar{\mu}}(U,V),$
$\mathbb{E}(\tau(x,V)-\tau(x,U))^{2}$ $\leq 4(\mathbb{E}d({\bar{\mu}},x)^{2})d_{\bar{\mu}}(U,V)^{2},$
$\text{and}\quad\big{|}\operatorname{\Sigma}(x,U,V)-\operatorname{\Sigma}(x,U,W)\big{|}$ $\leq 4(\mathbb{E}d({\bar{\mu}},x)^{2})^{\frac{1}{2}}d_{\bar{\mu}}(W,V).$
Proof.
Since $\lvert\tau(x,V)-\tau(x,U)\rvert\leq\lvert\langle\log_{\bar{\mu}}x,V\rangle_{\bar{\mu}}-\langle\log_{\bar{\mu}}x,U\rangle_{\bar{\mu}}\rvert+\lvert m(x,V)-m(x,U)\rvert$, the first estimate follows because $|m(x,V)-m(x,U)|\leq(\mathbb{E}d({\bar{\mu}},x))d_{\bar{\mu}}(U,V)$ and $|\langle X,V\rangle_{\bar{\mu}}-\langle X,U\rangle_{\bar{\mu}}|\leq d({\bar{\mu}},x)d_{\bar{\mu}}(U,V)$ when $X=\log_{\bar{\mu}}x$. Squaring this estimate, taking the expectation, and using the fact that $[\mathbb{E}d({\bar{\mu}},x)]^{2}\leq\mathbb{E}[d({\bar{\mu}},x)^{2}]$ produces the second inequality. The last estimate follows from the first two since $\|U\|=\|V\|=1$ and
$|\operatorname{\Sigma}(U,V)-\operatorname{\Sigma}(U,W)|$ $\leq(\mathbb{E}\tau(x,U)^{2}\mathbb{E}(\tau(x,V)-\tau(x,W))^{2})^{\frac{1}{2}}$
[…]
Proposition 3.2.
If $x$ is a random variable on $\mathcal{M}$ with $\mathbb{E}d({\bar{\mu}},x)<\infty$ then the centered random tangent field $V\mapsto\tau(x,V)$ induced by $x$ is almost surely Lipchitz-continuous and hence $\tau(x,\;\cdot\;)\in\mathcal{C}(S_{\bar{\mu}}\mathcal{M},\mathbb{R})$ almost surely.
Proof.
Proposition 3.1 yields the almost sure bound
$\lvert\tau(x,V)-\tau(x,U)\rvert\leq\bigl{(}\mathbb{E}d({\bar{\mu}},x)+d({\bar{\mu}},x)\bigr{)}d_{\bar{\mu}}(U,V).$
Thus $\tau(x,\;\cdot\;)$ is Lipchitz-continuous and $\tau(x,\;\cdot\;)\in\mathcal{C}(S_{\bar{\mu}}\mathcal{M},\mathbb{R})$ almost surely. ∎
3.2. Hölder-continuous Gaussian tangent fields
Corollary 3.3.
Let $\nu$ be a measure on $\mathcal{M}$ with $\int_{\mathcal{M}}d({\bar{\mu}},y)^{2}\nu(dy)<\infty$ and $\operatorname{\Sigma}$ the tangent covariance induced by $\nu$. Then there exists a version of the Gaussian tangent field induced by $\nu$ that is almost surely a Hölder continuous function from $S_{\bar{\mu}}\mathcal{M}\to\mathbb{R}$ for all Hölder exponents between $(0,1)$.
Proof.
Let $G$ be the Gaussian tangent field induced by $\nu$. By direct calculation,
$\mathbb{E}\bigl{(}G(U)-G(V)\bigr{)}^{2}$ $=\operatorname{\Sigma}(\nu,U,U)+\operatorname{\Sigma}(\nu,V,V)-2\operatorname{\Sigma}(\nu,U,V)=\mathbb{E}\bigl{(}\tau(x,U)-\tau(x,V)\bigr{)}^{2},$
where $x$ is $\mathcal{M}$-valued and distributed as $\mu$. By the middle estimate in Proposition 3.1,
$\mathbb{E}(G(U)-G(V))^{2}\leq K_{2}(\nu)d_{\bar{\mu}}(U,V)^{2}$
for some positive constant $K_{2}(\nu)$ depending on $\nu$. However, since $G(U)-G(V)$ is a centered Gaussian its higher moments can be bounded by powers of the second moment with the right homogeneity. Thus for any positive integer $p$ there exists a constant $K_{p}(\nu)$ so that
$\mathbb{E}|G(U)-G(V)|^{p}\leq K_{p}(\nu)d_{\bar{\mu}}(U,V)^{p}.$
With this estimate in hand, the result on the Hölder continuity follows from (1). ∎
Remark 3.4.
By Corollary 3.3, if $\gamma\in(0,1)$ then one can choose a $\gamma$-Hölder continuous version of a Gaussian tangent field which shares a tangent covariance with a given probability measure $\nu$ on $\mathcal{M}$ with $\int_{\mathcal{M}}d({\bar{\mu}},y)^{2}\nu(dy)<\infty$. Therefore, when discussing such Gaussian tangent fields, our convention is to consider only such Hölder continuous versions, without further comment.
3.3. Chaining and covering: proof of Theorem 1
The proof of Theorem 1 requires ideas from chaining and covering estimates. Optimal estimates are not the goal, given that simpler results suffice here. The proofs are standard, though the setting requires some adaptation. See [Kal97, Tal14] for more general and optimal results.
[…]
$N(\varepsilon)=\inf_{D\in\frac{\varepsilon}{2}\text{-cov}}|D|$
where $|D|$ is the cardinality of the set $D$.
Let $d\in(0,\infty)$ be such that
(3.2) $\limsup_{n\to\infty}\frac{1}{n}\log_{2}N(2^{-n})\leq d.$
When $\mathcal{M}$ is some Euclidean space $\mathbb{R}^{k+1}$ write $N_{k}(\varepsilon)$ for the covering number of the $k$-dimensional unit sphere, and let $d_{k}\in(0,\infty)$ be such that
(3.3) $\limsup_{n\to\infty}\frac{1}{n}\log_{2}N_{k}(2^{-n})\leq d_{k}.$
Lemma 3.5.
For any stratified $\operatorname{CAT}(\kappa)$ space $\mathcal{M}$, the number $d$ in (3.2) can be chosen to be finite.
Proof.
Suppose that $\{R_{i}\}_{i=1}^{\ell}$ is the list of strata of $\mathcal{M}$ and $r_{i}$ is the dimension of $R_{i}$. Suppose that $\mathcal{B}_{i}(\varepsilon)$ is an $\varepsilon$-cover of $R_{i}\cap S_{\bar{\mu}}\mathcal{M}$. Then $\bigcup_{i=1}^{n}\mathcal{B}_{i}(\varepsilon)$ is an $\varepsilon$-cover of $S_{\bar{\mu}}\mathcal{M}$. Hence $d$ can be chosen such that $d\leq\sum_{i=1}^{\ell}d_{r_{i}}$, which is finite. ∎
Proof of Theorem 1.
The arguments loosely follow those in [Kal97, Theorem 3.23].
Fix a sufficiently large, positive integer $M$ and let $d$ be the constant from (3.3). For each $n\geq M$, let $\widehat{D}_{n}$ be the centers of the balls in a $2^{-(n+1)}$-cover such that $|\widehat{D}_{n}|\leq\frac{3}{2}N(2^{-(n+1)})\leq 2\cdot 2^{d(n+1)}$. By definition of $N(2^{-(n+1)})$ and $d$ from (3.2) and (3.3), this is always possible for $M$ large enough. Now define $D_{n}=\bigcup_{k=M}^{n}\widehat{D}_{n}$ and observe that $D_{n}\subset D_{n+1}$, and $|D_{n}|<\mathcal{N}_{n}=\sum_{k=M}^{n}2\cdot 2^{d(n+1)}\leq K2^{dn}$ for some $K$ depending on $d$ but not $n$. Set
$\xi_{k}=\sup\bigl{\{}|f(V)-f(U)|\,\bigm{|}\,U,V\in D_{k}\text{ and }d_{\bar{\mu}}(U,V)\leq 2^{-k}\bigr{\}}.$
For any integer $a>\max(d,1)$, observe that
$\xi_{k}^{a}\ \leq\sum_{\begin{subarray}{c}U,V\in D_{k}\\ d_{\bar{\mu}}(U,V)\leq 2^{-k}\end{subarray}}|f(V)-f(U)|^{a}\,.$
By assumption $\mathbb{E}|f(V)-f(U)|^{a}\leq C_{a}d_{\bar{\mu}}(U,V)^{a}$ for some positive constant $C_{a}$ independent of $U$ and $V$. Since |$\bigl{|}\bigl{\{}U,V\in D_{k}\,\bigm{|}\,d_{\bar{\mu}}(U,V)\leq 2^{-k}\bigr{\}}\bigr{|}\leq K2^{dk}$,
$\mathbb{E}\xi_{k}^{a}\leq K2^{-ka}2^{dk}=K2^{-k(a-d)}$
for some possibly different constant $K$. Therefore, for any $\gamma\in(0,(a-d)/a)$ and $a>d$,
$\sum_{k}\mathbb{E}\xi_{k}^{a}2^{\gamma ka}\leq C\sum_{k}2^{-k(a-a\gamma-d)}.$
Since $a-a\gamma-d>0$ this sum converges. Hence for some random but finite $K_{0}$, $\sum_{k}\xi_{k}^{a}2^{\gamma ka}\leq K_{0}$ almost surely. Consequently, for some random constant $K$ and any $m\geq M$,
$\sup\bigl{\{}|f(U)-f(V)|\,\bigm{|}\,U,V\in\bigcup_{k}D_{k},d_{\bar{\mu}}(V,U)\leq 2^{-m}\bigr{\}}\leq\Bigl{(}\sum_{k\geq m}\xi_{k}^{a}\Bigr{)}^{\frac{1}{a}}\leq K2^{-m\gamma}$
almost surely, so $f$ is almost surely Hölder continuous with exponent $\gamma$ on the set $\bigcup_{k}D_{k}$. Picking $a$ large enough makes $(a-d)/a$ arbitrarily close to $1$, so $\gamma$ can be chosen arbitrarily close to $1$.
Since $\bigcup_{k}D_{k}$ is dense in $S_{\bar{\mu}}\mathcal{M}$, a Hölder continuous version $\hat{f}$ of $f$ can be defined by setting $\hat{f}(U)=f(U)$ for $U\in\bigcup_{k}D_{k}$ and otherwise setting $\hat{f}(U)=\lim f(U_{k})$ for any sequence $U_{k}\in\bigcup_{k}D_{k}$ with $U_{k}\rightarrow U$. Since $\hat{f}(U_{k})=f(U_{k})$ is continuous on $\bigcup_{k}D_{k}$ this definition is independent of the chosen sequence $U_{k}\rightarrow U$. The resulting $\hat{f}$ is Hölder continuous for $\gamma<1$.
To see that $\hat{f}$ is a version of the field $f$, (3.1) implies that $f(U_{k})$ converges to $f(U)$ in probability while $\hat{f}(U_{k})$ converges to $\hat{f}(U)$ by continuity of $\hat{f}$ inherent in its construction. Hence $\mathbf{P}\bigr{(}f(U)=\hat{f}(U)\bigr{)}=1$ for all $U\in S_{\bar{\mu}}\mathcal{M}$ since $f(U_{k})=\hat{f}(U_{k})$. ∎
4. CLT for random tangent fields
Definition 4.1.
Fix a measure $\mu$ on a $\operatorname{CAT}(\kappa)$ space $\mathcal{M}$ and a sequence $\mathbf{x}=(x_{1},x_{2},\ldots)$ in $\mathcal{M}^{\mathbb{N}}$ of mutually independent random variables $x_{i}\sim\mu$. Denote by $g_{i}(V)=\tau(x_{i},V)$ the centered random tangent field induced by $\mu$ (Definition 2.5), where ${\bar{\mu}}$ there is the Fréchet mean of $\mu$. Then
$G_{n}(V)$ $=\frac{1}{\sqrt{n}}\sum_{k=1}^{n}g_{i}(V)$
is the empirical tangent field induced by $\mu$. In contrast, the average of $g_{1},\dots,g_{n}$ is
${\overline{g}}_{n}(V)$ $=\frac{1}{n}\sum_{k=1}^{n}g_{i}(V)=\frac{1}{\sqrt{n}}G_{n}.$
The $g_{i}$ are centered, so their expectations vanish. Their averages ${\overline{g}}_{n}(V)$ have the classical Law of Large Numbers scaling and hence converge to 00, while the $G_{n}$ have the classical CLT scaling selected to stablize the fluctuations. The following central limit theorem for the tangent field $G_{n}$ makes this precise, intrinsically capturing the asymptotics of sample means in the linear, albeit nontraditional setting of Gaussian fields. The proof occupies Section 5.
Theorem 2 (Random tangent field CLT).
Fix a localized measure $\mu$ on a stratified $\operatorname{CAT}(\kappa)$ space $\mathcal{M}$. Let $G$ be the Gaussian tangent field induced by $\mu$ (Definition 2.8), and let $G_{n}=\sqrt{n}\,{\overline{g}}_{n}$ be the empirical tangent fields (Definition 4.1) for the collection of independent random variables $x_{i}$ each distributed as $\mu$. Then the $G_{n}$ converge to $G$
[…]
in distribution as $n\rightarrow\infty$.
The following asserts that the limiting Gaussian tangent field $G$ and the empirical tangent fields $G_{n}$ can be constructed on the same probability space so that the convergence is almost sure as functions in $\mathcal{C}(S_{\bar{\mu}}\mathcal{M},\mathbb{R})$. Its proof is at the end of Section 5.
Corollary 4.2.
In the setting of Theorem 2, there exist versions of $G_{n}$ and $G$ so that $G_{n}\to G$ almost surely in $\mathcal{C}(S_{\bar{\mu}}\mathcal{M},\mathbb{R})$ equipped with the supremum norm and so that $G$ is almost surely Hölder continuous for any Hölder exponent less than $1$.
5. Proof of the CLT for random tangent fields
This subsection culminates in the proofs of Theorem 2 and Corollary 4.2. It begins by collecting a number of auxiliary results and observations.
Throughout the subsection, assume the situation of Definition 4.1. For notational convenience in this setting, write
$X_{i}=\log_{\bar{\mu}}x_{i},\qquad\text{so}\qquad X_{i}\sim{\widehat{\mu}}=(\log_{\bar{\mu}})_{\sharp}\mu$
as in Definition 1.6. Also use $\|V\|=d({\bar{\mu}},\exp_{\bar{\mu}}V)$ to denote the length of a vector $V\in T_{\bar{\mu}}\mathcal{M}$. Lastly, write $\operatorname{\Sigma}(V,U)$ for the covariance $\operatorname{\Sigma}(\mu,V,U)$ of $\mu$ and $F=F_{\mu}$ for the Fréchet function of $\mu$ in (1), as usual.
[…]
Lemma 5.1.
The Gaussian random tangent field $G$ can be understood as a random variable taking values in $L^{2}(S_{\bar{\mu}}\mathcal{M},\mathbb{R})$ with finite second moment.
Proof.
As the angular pairing at ${\bar{\mu}}$ is continuous [MMT23a, Lemma 1.21] so is the covariance $\operatorname{\Sigma}(V,V)$. Since $S_{\bar{\mu}}\mathcal{M}$ is compact, we know that $\operatorname{\Sigma}(V,V)$ is bounded on $S_{\bar{\mu}}\mathcal{M}$ and in particular
[…]
Lemma 5.2.
For each $n\geq 1$, the empirical tangent field $G_{n}$ is almost surely Lipchitz-continuous. In particular, $G_{n}\in\mathcal{C}(S_{\bar{\mu}}\mathcal{M},\mathbb{R})$ with probability $1$.
Proof.
There is an almost sure bound
$\lvert G_{n}(V)-G_{n}(U)\rvert\leq\Bigl{(}\sqrt{n}\mathbb{E}\lVert X_{1}\rVert+\frac{1}{\sqrt{n}}\sum_{i=1}^{n}\lVert X_{i}\rVert\Bigr{)}d_{\bar{\mu}}(U,V)$
since $|m(x_{i},\!V)-m(x_{i},\!U)|\!\leq\!(\mathbb{E}\lVert X_{i}\rVert)d_{\bar{\mu}}(U,\!V)$ and $|\langle X_{i},\!V\rangle_{\bar{\mu}}-\langle X_{i},\!U\rangle_{\bar{\mu}}|\!\leq\!\lVert X_{i}\rVert d_{\bar{\mu}}(U,\!V)$. Thus $G_{n}$ is Lipchitz-continuous and $G_{n}\in\mathcal{C}(S_{\bar{\mu}}\mathcal{M},\mathbb{R})$ almost surely. ∎
Lemma 5.3.
For any $p\in\mathbb{N}$, there exists a universal constant $C_{p}$ so that
$\sup_{n\geq 1}\,\mathbb{E}\>|G_{n}(V)|^{2p}\leq C_{2p}{\gamma_{2p}}$
where $\gamma_{2p}=\mathbb{E}\>\lVert X_{1}\rVert^{2p}$.
Proof.
The bound for $p>2$ follows by reasoning as in the $p=2$ case, which follows from
$\mathbb{E}\>\lvert G_{n}(V)\rvert^{4}$ $=\frac{1}{n^{2}}\sum_{\begin{subarray}{c}i,j=1\\ i\neq j\end{subarray}}^{n}\mathbb{E}g_{i}(V)^{2}\mathbb{E}g_{j}(V)^{2}+\frac{1}{n^{2}}\sum_{i=1}^{n}\mathbb{E}g_{i}(V)^{4}$
$\leq\frac{n(n-1)\gamma_{2}^{2}}{n^{2}}+\frac{n\gamma_{4}}{n^{2}}$
$\leq C_{4}\gamma_{4}.\qed$
Let $\mathcal{F}_{n}$ be the $\sigma$-algebra generated by $G_{j}(V)$ such that $j\leq n$ and $V\in T_{\bar{\mu}}\mathcal{M}$.
Proposition 5.4.
For fixed $V\in S_{\bar{\mu}}\mathcal{M}$, the assignment $n\mapsto G_{n}(V)$ is a martingale relative to the filtration $\mathcal{F}_{n}$. In fact, any fixed $V_{i}\in T_{\bar{\mu}}\mathcal{M}$ and $c_{i}\in\mathbb{R}$ yield a martingale
$n\mapsto\sum_{i=1}^{k}c_{i}G_{n}(V_{i}).$
Proof.
By construction $G_{n}(V)$ is adapted to $\mathcal{F}_{n}$ and $\sup_{n}\mathbb{E}|G_{n}(V)|<\infty$ by the estimates in Lemma 5.3. $\mathbb{E}(G_{n+k}(V)\mid\mathcal{F}_{n})=G_{n}(V)$ because each $g_{i}(V)$ has mean zero and is independent of the random variables $\{g_{k}(V)\mid k\neq i\}$. The claim about finite linear combinations follows by the same reasoning. For more on the definition of a martingale and further details see [Wil91]. ∎
Corollary 5.5.
For any $V\in S_{\bar{\mu}}\mathcal{M}$,
$\mathbf{P}\Bigl{(}\sup_{n}\lvert G_{n}(V)\rvert>\beta\Bigr{)}\leq\frac{C_{2p}\gamma_{2p}}{\beta^{2p}}.$
Proof.
The estimate follows from the classical Doob martingale bound and the moment bounds in Lemma 5.3; see [Dur19, Section 4.4] or [Wil91], for example. ∎
Lemma 5.6.
Given $p\in\mathbb{N}$ there is a universal constant $C_{p}$ so that for any $V,U\!\in\nolinebreak\!S_{\bar{\mu}}\mathcal{M}$,
$\mathbb{E}\>\lvert G_{n}(V)-G_{n}(U)\rvert^{2p}\leq C_{2p}{(1+\gamma_{2p})}d_{\bar{\mu}}(U,V)^{2p},$
where $\gamma_{2p}=\mathbb{E}\>\|X_{1}\|^{2p}$. Additionally, for any $V,U\in S_{\bar{\mu}}\mathcal{M}$,
$\mathbf{P}\Bigl{(}\sup_{n}\lvert G_{n}(V)-G_{n}(U)\rvert>\beta\Bigr{)}\leq\frac{C_{2p}(1+\gamma_{2p})}{\beta^{2p}}d_{\bar{\mu}}(U,V)^{2p}$
for any $\beta>0$.
Proof.
As before, the proofs for higher moments follow the same pattern as the estimate for $p=2$, which is proved as follows. Begin by observing that
$\mathbb{E}\>\lvert g_{i}(V)-g_{i}(U)\rvert^{2}$ $\leq 2\lvert m(V)-m(U)\rvert^{2}+2\mathbb{E}\>\lvert\langle\log X_{i},V\rangle_{\bar{\mu}}-\langle\log X_{i},U\rangle_{\bar{\mu}}\rvert^{2}$
$\leq 2\lvert m(V)-m(U)\rvert^{2}+2\mathbb{E}\>\lvert\log X_{i}\rvert^{2}d_{\bar{\mu}}(V,U)^{2}$
$\leq 2\bigl{(}1+\mathbb{E}\>\lvert\log X_{i}\rvert^{2}\bigr{)}d_{\bar{\mu}}(V,U)^{2},$
so
$\mathbb{E}\>\lvert g_{i}(V)-g_{i}(U)\rvert^{4}\leq C\bigl{(}1+\mathbb{E}\>\lvert\log X_{i}\rvert^{4}\bigr{)}d_{\bar{\mu}}(V,U)^{4}.$
Again, using the independent increments produces
$\mathbb{E}\>\lvert G_{n}(V)-G_{n}(U)\rvert^{4}$ $\leq\frac{1}{n^{2}}\!\sum_{\begin{subarray}{c}j,i=1\\ i\neq j\end{subarray}}^{n}\mathbb{E}\>|g_{i}(V)-g_{i}(U)|^{2}\,\mathbb{E}\>|g_{j}(V)-g_{j}(U)|^{2}$
$\phantom{\leq}\ +\frac{1}{n^{2}}\!\sum_{i=1}^{n}\mathbb{E}\>|g_{i}(V)-g_{i}(U)|^{4}$
(5.1) $\leq C(1+\gamma_{4})d_{\bar{\mu}}(V,U)^{4}.$
To get the final estimate from (5.1), apply Doob’s Martingale inequality, as in the proof of Corollary 5.5, to $n\mapsto G_{n}(V)-G_{n}(U)$, which is a martingale by Proposition 5.4. ∎
5.2. Convergence
Begin with the following result on finite-dimensional distributions.
Lemma 5.7.
For any $k\in\mathbb{N}$ and vectors $V_{1},\dots,V_{k}\in S_{\bar{\mu}}\mathcal{M}$ in the unit tangent sphere, the vector $\xi_{n}=\bigl{(}G_{n}(V_{1}),\dots,G_{n}(V_{k})\bigr{)}\in\mathbb{R}^{m}$ converges in distribution to the Gaussian vector $\xi=\bigl{(}G(V_{1}),\dots,G(V_{k})\bigr{)}$ with covariance matrix $\operatorname{\Sigma}_{ij}=\{\operatorname{\Sigma}(V_{i},V_{j})\mid 1\leq i,j\leq k\}$.
Proof.
The increments $\xi_{n}-\xi_{n-1}$ are independent, so the result follows from the standard central limit theorem for independent random variables on $\mathbb{R}^{m}$ after noting that the $\xi_{n}$ share the same covariance matrix $\operatorname{\Sigma}$. ∎
The next lemma uses the notion of tightness (see [Bil13, Section 5.1] for example): a family of measures on a metric space $S$ with the Borel $\sigma$-field $\mathcal{S}$ is tight if for every $\varepsilon>0$ some compact set $K\subseteq S$ has $\nu(K)>1-\varepsilon$ for every measure $\nu$ in the family. The interest in tightness stems from the fact that by Prokhorov’s Theorem a tight sequence of probability measures is relatively compact in distribution and hence must have a subsequence which converges in distribution.
When considering convergence in distribution on a space of continuous functions (such as $\mathcal{C}(S_{\bar{\mu}}\mathcal{M},\mathbb{R})$, where $G_{n}$ lies by Lemma 5.2), the Arzelà–Ascoli Theorem can prove tightness by controlling the modulus of continuity. If $h\in\mathcal{C}(S_{\bar{\mu}}\mathcal{M},\mathbb{R})$ and
$w(h,r)=\sup\bigl{\{}|h(V)-h(U)|\,\bigm{|}\,U,V\in S_{\bar{\mu}}\mathcal{M}\text{ and }d_{\bar{\mu}}(V,U)\leq r\bigr{\}},$
$K(c_{0},c_{1},r_{0})=\bigl{\{}h\in\mathcal{C}(S_{\bar{\mu}}\mathcal{M},\mathbb{R})\mid|h(V)|\leq c_{0}\text{ for some }V\in S_{\bar{\mu}}\mathcal{M}\text{ and }w(h,r_{0})\leq c_{1}\bigr{\}}$
is relatively compact for any positive $c_{0}$, $c_{1}$, and $r_{0}$. By construction, the collection of functions in $K(c_{0},c_{1},r_{0})$ is equicontinuous, so any limit point in the supremum topology is also continuous.
Lemma 5.8.
If $\mathcal{M}$ is stratified then the sequence $\{G_{n}\}$ is tight in $\mathcal{C}(S_{\bar{\mu}}\mathcal{M},\mathbb{R})$.
Proof.
The arguments are similar to [Kal97, Theorem 16.9] and more generally [Tal14]. Fix $\mathcal{M}$ and the collection $\{D_{n}\}_{n=M}^{\infty}$ of covering ball centers as for Theorem 1, and set
$\xi_{n,k}=\sup\bigl{\{}|G_{n}(V)-G_{n}(U)|\,\bigm{|}\,U,V\in D_{k}\text{ and }d_{\bar{\mu}}(U,V)\leq 2^{-k}\bigr{\}}.$
For any even integer $a>\max(d,1)$ where $d$ is the constant from (3.3), observe that
$\xi_{n,k}^{a}\leq\sum_{\begin{subarray}{c}U,V\in D_{k}\\ d_{\bar{\mu}}(U,V)\leq 2^{-k}\end{subarray}}|G_{n}(V)-G_{n}(U)|^{a}$
and, by Lemma 5.6, $\mathbb{E}\>|G_{n}(V)-G_{n}(U)|^{a}\leq C_{a}d_{\bar{\mu}}(U,V)^{a}$ for some positive constant $C_{a}$ independent of $U$ and $V$. Since the set $\bigl{\{}U,V\in D_{k}\,\bigm{|}\,d_{\bar{\mu}}(U,V)\leq 2^{-k}\bigr{\}}$ has size $\leq K2^{dk}$,
(5.2) $\mathbb{E}\>\xi_{n,k}^{a}\leq K2^{-ka}2^{dk}=K2^{-k(a-d)}$
for some possibly different constant $K$. Next, given any $U,V\in S_{\bar{\mu}}\mathcal{M}$ that satisfies $d_{\bar{\mu}}(U,V)\leq 2^{-m}$ there exist two sequences $\{U_{j}\}_{j=m}^{\infty}$ and $\{V_{j}\}_{j=m}^{\infty}$ with $U_{k},V_{k}\in D_{k}$ and $\lim U_{k}=U$ and $\lim_{k}V_{k}=V$. Then
$|G_{n}(U)-G_{n}(V)|\leq|G_{n}(U_{m})-G_{n}(V_{m})|$ $+\sum_{k=m}^{\infty}|G_{n}(U_{m})-G_{n}(U_{m+1})|$
$+\sum_{k=m}^{\infty}|G_{n}(V_{m})-G_{n}(V_{m+1})|$
since $G_{n}$ is continuous. This in turn implies that $|G_{n}(U)-G_{n}(V)|\leq C\sum_{k\geq m}\xi_{n,k}$ for some universal positive $C$, and hence $w(h,2^{-m})\leq C\sum_{k\geq m}\xi_{n,k}$. Since $a>1$, applying the triangle inequality to this estimate and then (5.2) produces
$\mathbb{E}\bigl{(}w(G_{n},2^{-m})^{a}\bigr{)}\leq\Bigl{(}C\sum_{k\geq m}\bigl{(}\mathbb{E}\xi_{n,k}^{a}\bigr{)}^{\frac{1}{a}}\Bigr{)}^{a}\leq C\Bigl{(}\sum_{k\geq m}2^{-k\bigl{(}1-\frac{d}{a}\bigr{)}}\Bigr{)}^{a}\leq C2^{-m(a-d)}.$
Since $a-d>0$, this implies that
$\lim_{m\to\infty}\limsup_{n\to\infty}\mathbb{E}\bigl{[}w(G_{n},2^{-m})\wedge 1\bigr{]}=0,$
which shows the tightness of $\{G_{n}\}$. ∎
Lemma 5.9.
If $\mathcal{M}$ is stratified then the sequence $G_{n}$ converges in distribution to $G$.
Proof.
The result follows from [Bil13, Theorem 7.5] or [Kal97, Theorem 16.5] in the setting of continuous functions on $\mathbb{R}^{d}$. We sketch the proof for completeness and to translate it to our setting on a stratified space $\mathcal{M}$ rather than $\mathbb{R}^{d}$. The result follows from the classical result that $G_{n}$ converges in distribution to $G$ if and only if every subsequence in $\{{\widehat{G}}_{k}=G_{n_{k}}\}_{k}$ contains a further subsequence which $\{{\widehat{G}}_{k_{j}}\}$ which converges in distribution to $G$ (see [Bil13, Theorem 2.6], for example). Since the $\{G_{n}\}$ are tight by Lemma 5.8, Prokhorov’s Theorem in the setting of continuous functions [Bil13, Theorem 7.1] implies that any subsequence $\{{\widehat{G}}_{k}=G_{n_{k}}\}_{k}$ contains a further subsequence $\{{\widehat{G}}_{k_{j}}\}$ which converges in distribution to a limit which we denote by ${\widehat{G}}\in\mathcal{C}(S_{\bar{\mu}}\mathcal{M},\mathbb{R})$. Since ${\widehat{G}}$ is the uniform limit of continuous functions it is also continuous. Since both ${\widehat{G}}$ and $G$ are continuous, they are determined completely by their finite-dimensional distributions. However, since $\bigl{(}{\widehat{G}}_{k_{j}}(V_{1}),\dots,{\widehat{G}}_{k_{j}}(V_{n})\bigr{)}$ converges to $\bigl{(}G(V_{1}),\dots,G(V_{n})\bigr{)}$ for any collection $V_{1},\dots,V_{n}\in S_{\bar{\mu}}\mathcal{M}$, it must be that $G$ and ${\widehat{G}}$ have the same distribution. ∎
Proof of Corollary 4.2.
$\mathcal{C}(S_{\bar{\mu}}\mathcal{M},\mathbb{R})$ is a complete separable metric space when endowed with the supremum norm. Convergence then follows from [Kal97, Theorem 3.30]. As each $G_{n}$ is continuous and $G$ is their uniform limit, $G$ is almost-sure continuous. ∎